the length of one side a triangular lot is 6m less than three times the length of the second side. the third side is 8m longer than the first side. the perimeter of the lot is 80 m. Find the length of all three sides
Answer by skartikey(21)

let sides are a, b, c
a=3b-6
c=a+8
a+b+c=80
3b-6+b+a+8=80
3b-6+b+3b-6+8=80
7b-4=80
7b=84
b=12
a=3b-6=3(12)-6=30
c=a+8=30+8=38